OpenStudy (anonymous):

The equation of the circle whose center is (0, 0) and radius is length 9 is? 3,9, or 81.

OpenStudy (anonymous):

x^2 + y^2 = r^2, general form So if r = 9 Then equation is: x^2+y^2=81